What is the distance between the points (21, -30) and (3, 8)?
Answer by jim_thompson5910(35256)

d = sqrt((x2-x1)^2+(y2-y1)^2)


d = sqrt((3-21)^2+( 8-( -30))^2)


d = sqrt((3-21)^2+( 8-( -30))^2)


d = sqrt((-18)^2+(38)^2)


d = sqrt(324+1444)


d = sqrt(1768)


d = 2*sqrt(442)


So exact distance between the two points is 2*sqrt(442) units.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------